# Advances and Open Challenges in Federated Learning with Foundation Models

### v2

@@ Abstract
The integration of Foundation Models (FMs) with Federated Learning (FL) presents a transformative paradigm in Artificial Intelligence (AI), offering enhanced capabilities while addressing concerns of privacy, data decentralization, and computational efficiency. This paper provides a comprehensive survey of the emerging field of Federated Foundation Models (FedFM), elucidating their synergistic relationship and exploring novel methodologies, challenges, and future directions that the FL research field needs to focus on in order to thrive in the age of foundation models. A systematic multi-tiered taxonomy is proposed, categorizing existing FedFM approaches for model training, aggregation, trustworthiness, and incentivization. Key challenges, including how to enable FL to deal with high complexity of computational demands, privacy considerations, contribution evaluation, and communication efficiency, are thoroughly discussed. Moreover, the paper explores the intricate challenges of communication, scalability and security inherent in training/fine-tuning FMs via FL, highlighting the potential of quantum computing to revolutionize the training, inference, optimization and data encryption processes. This survey underscores the importance of further research to propel innovation in FedFM, emphasizing the need for developing trustworthy solutions. It serves as a foundational guide for researchers and practitioners interested in contributing to this interdisciplinary and rapidly advancing field.
@@
Federated learning, foundation models, large language models, efficient training/aggregation, trustworthiness, incentivization, contribution evaluation, quantum computing.
@@ I Introduction
The Artificial Intelligence (AI) landscape is undergoing significant advancement propelled by the emergence of Foundation Models (FMs) [1, 2]. These models are named for their essential role as the foundational backbone for a wide range of learning tasks with diverse modalities [3]. While technically representing an evolutionary rather than revolutionary development in AI [2], FMs have significantly reshaped model training and deployment practices through their widespread application. FMs are typically pre-trained on large-scale datasets using self-supervised learning methods. This pre-training equips them with general and adaptable knowledge, applicable across diverse tasks. Such practice arises from the high costs associated with FM training and the relatively limited scale of available datasets for these tasks, rendering the traditional training-from-scratch paradigm impractical.
Recently, FMs with their diverse designs and learning processes have achieved remarkable success in tackling complex tasks that were once considered too difficult or even impossible. Particularly in Natural Language Processing (NLP), FMs such as those in the GPT series [4, 5, 6], LLaMa [7] and PaLM [8] have demonstrated extraordinary capabilities. In computer vision, Segment Anything [9] has emerged as a standout performer. In the realm of Generative AI models, Stable Diffusion [10] has garnered significant attention. Notably, DALL-E [11] and CLIP [12] have demonstrated outstanding performance in multi-modal tasks, bridging the understanding of textual and visual data. Nevertheless, the success of FMs comes at a high cost. The enhanced capabilities necessitate a substantial amount of high-quality training data and computational resources. This underscores the potential value of a collaborative learning paradigm, one that enables the distribution of both data resources and computational loads.
Federated learning (FL) [13], a collaborative learning approach, has received significant interest for its potential in enhancing privacy, leveraging distributed datasets, and distributing computational loads. Training/fine-tuning FMs via FL offers several advantages. FL can broaden the data horizons for FMs by aggregating diverse data sources from both the public and private sectors [14], including research institutions and industries. It can also distribute computational loads among multiple participants, leading to more efficient resource utilization and mitigating the risk of AI monopolies by major technology companies. On the other hand, the robustness of FMs strengthens the effectiveness of FL, particularly in managing non-iid (non-independent and identically distributed) data. Moreover, the flexibility of FMs in adapting to various downstream tasks facilitates easier personalization of FL models. These benefits highlight the promising potential of integrating FMs with FL.
@@
The paper is organized as follows. Section II introduces FMs and FL, concepts relevant to our discussion. Section III describes the FedFM taxonomy, which encompasses several key aspects: training and aggregation methods, ensuring trustworthiness, creating effective incentive mechanisms, and managing evaluation. Section IV focuses on aggregation, communication and computational efficiency, highlighting the nuanced considerations required for the effective deployment of FedFM. Section V addresses the critical areas of Byzantine-robustness and privacy for FedFM. Section VI highlights the complexity of developing effective incentive mechanisms for FedFM. Section VII describes and compares prevailing FedFM frameworks. Section VIII discusses open research challenges of FedFM in terms of memory, computation, and communication, while also delving into potential future directions involving quantum computing techniques and effective evaluation. Finally, section IX concludes the paper.
II Terminology
This section aims to provide an overview of FMs and FL, outlining the fundamental concepts and historical development trajectories that have propelled their current prominence in the field.
@@ II-A Foundation Models
The term FMs was coined in [2] to denote the role of such models as a foundational base, from which numerous task-specific models can be built through adaptation. FMs emphasize the stability of their architecture during training and the consistency of parameters throughout the adaptation phase. This term also signifies a profound shift in AI research and deployment, which began in the field of NLP, where the synergy of large-scale models and transfer learning [20] resulted in the development of powerful models. The scale of FMs manifests in three key dimensions: 1) the sizes of the models, 2) the extensive computational resources required for training, and 3) the substantial volume of data consumed. For example, GPT-2 [4] consists of 1.3 billion parameters with a memory footprint of approximately 2.6 GB, while GPT-3 [5] consists of 175 billion parameters with a memory footprint of around 350 GB in half-precision format. GPT-4 [6] further extends this scale to over one trillion parameters. In terms of computational resources, training FMs demands substantial power. For instance, the LLaMA [7] model, requires 2,048 NVIDIA A100 GPUs for a duration of 21 days for training, while the vision transformer in CLIP [21] requires 8 core TPUv3 for approximately 30 days for training. Additionally, FMs consume vast amounts of training data. GPT-3 [5], for instance, is trained on a dataset with 300 billion tokens, equivalent to hundreds of gigabytes of data. Research suggests a potential depletion of high-quality language data by 2026 [22], raising concerns that data availability might eventually impede the advancement of FMs.
@@ II-B Federated Learning
FL aims to collaboratively train models across multiple decentralized data owners holding potentially sensitive local data in a privacy-preserving manner [13]. The essence of FL lies in its ability to learn a shared model by aggregating locally computed updates, rather than directly accessing or sharing the raw data. This machine learning (ML) paradigm not only enhances privacy and security, but also enables the utilization of distributedly owned data for model training, making it particularly suitable for mission critical applications, such as healthcare [23] and finance [24], where data privacy is of paramount concern.
@@ III-A Efficient Federated Training and Aggregation: The Foundation of FedFM
This part of the review underscores the evolving landscape of on FedFM research, detailing the challenges, strategies, and innovations in this interdisciplinary field. The focus on scaling, communication, and computational efficiency reflects the nuanced considerations required to effectively deploy FedFM.
@@ III-B Trustworthiness: A Crucial Pillar of FedFM Integrity
Trustworthy FedFM address the critical areas of Byzantine-robustness and privacy preservation in the context of FedFM, exploring the challenges and proposing solutions to ensure the integrity and security of these systems. We summarize key trends in each of these areas:
  • Byzantine Robustness: Research in this area discusses how poisoning attacks, both untargeted and targeted, aim to compromise the integrity of global models in FL systems. Untargeted attacks disrupt the training process to prevent convergence, while targeted attacks manipulate the model output subtly. The complexity and heterogeneity of FedFM training tasks make these attacks particularly challenging to execute and defend against. Byzantine-robust aggregation rules are also being proposed to counteract poisoning attacks, including geometrical outlier detection, top performance selection, and other hybrid schemes. However, the effectiveness of these defenses has been questioned in FedFM settings, calling for novel approaches tailored to its complexities.
@@
This part of the review on FedFM research underscores the intricate balance required to achieve trustworthiness, emphasizing the need for innovative solutions to address the dual challenges of Byzantine-robustness and privacy preservation. The complexity of FMs and the FL environment necessitates a re-evaluation of traditional defense mechanisms and development of new strategies to ensure the integrity and privacy of such advanced learning systems.
@@ IV Towards Efficient FedFM
In this section, we review and discuss current literature on the topic of efficient collaborative model training/fine-tuning and aggregation for FedFM, focusing on emerging settings and scenarios that arise as a result of combining FL with FMs.
@@
14:        $\bm{\theta}_{i}^{t}\leftarrow\bm{\theta}^{t+1}-\eta\nabla\mathcal{L}(\bm{\theta}^{t};b)$
15:        How to improve the computational efficiency so that FMs can be efficiently hosted and updated on FL clients?
@@
18:  return $\bm{\theta}_{i}^{t}$ to the server How to improve communication efficiency when $\bm{\theta}_{i}^{t}$ is large?
19:
20:  procedure Aggregate($\{\bm{\theta}_{i}^{t+1}\}_{i=1}^{m}$)
21:  FedAvg [25]: $\bm{\theta}^{t+1}\leftarrow\frac{1}{m}\sum_{i=1}^{m}\bm{\theta}_{i}^{t+1}$
22:  or
23:  Aggregation methods for FMs to enhance performance?
@@ IV-A FedFM Training and Aggregation
To grasp the challenges involved in integrating FMs into the FL process, a revision of the general FL procedure is necessary. FL typically operates under the control and decision-making of the central server with the primary purpose of collaboratively training a Deep Learning (DL) model $\bm{\theta}$ across $N$ clients over multiple rounds, as outlined in Algorithm 1. The integration of FMs, with their considerable sizes, introduces challenges at various stages of the FL process. These challenges can be distilled into three critical research questions (highlighted in blue in Algorithm 1):
  1. 1.
@@
Therefore, our survey examines innovative methods and strategies useful for FedFM from three dimensions: 1) aggregation methods for enhanced effectiveness, 2) techniques for boosting computational efficiency, and 3) techniques for improving communication efficiency. We further break down each dimension into sub-areas for more detailed analysis (Fig. 2).
@@ IV-B Scaling Efforts for FedFM
TABLE III: Sizes of FMs under FL Setting
Methods Year $\downarrow$ Size (M)
FedPAQ [30] 2020 0.2
HeteroFL [31] 2020 11
LotteryFL [32] 2020 138
FjORD [33] 2021 11
H-FL [34] 2021 138
PruneFL [35] 2022 132
FedPM [36] 2022 12
FedTiny [37] 2022 132
SoteriaFL [38] 2022 0.05
FedPrompt [39] 2022 223
FedBERT [40] 2022 117
FedCLIP [41] 2023 85
FedPEFT [42] 2023 85
FedPETuning [43] 2023 125
SLoRA [44] 2023 67
FedOBD [45] 2023 17
FedIT [46] 2023 7,000
FwdLLM [47] 2024 7,000
We list the methods reviewed alongside with the largest base models involved in their evaluations in Table III and Table IV. The substantial difference for model sizes between existing FL research and centralized FMs might cast doubts on the feasibility and practicality of the FL techniques for FedFM, calling for further experimental studies involving large-scale FMs. Such an effort depends on the availability of frameworks that can streamline implementation. Thus, in Section VII, we analyze the support for FMs in existing FL frameworks.
@@
Methods Year $\downarrow$ Size (M)
BERT[48] 2018 340
GPT1 [49] 2018 117
GPT2 [4] 2019 1,500
GPT3 [5] 2020 175,000
ViT [50] 2020 632
LLAMA [7] 2023 70,000
LLAVA [51] 2024 15,000

### v4

# Advances and Open Challenges in Federated Foundation Models

@@ Abstract
The integration of Foundation Models (FMs) with Federated Learning (FL) presents a transformative paradigm in Artificial Intelligence (AI). This integration offers enhanced capabilities, while addressing concerns of privacy, data decentralization and computational efficiency. This paper provides a comprehensive survey of the emerging field of Federated Foundation Models (FedFM), elucidating their synergistic relationship and exploring novel methodologies, challenges, and future directions that the FL research field needs to focus on in order to thrive in the age of FMs. A systematic multi-tiered taxonomy is proposed, categorizing existing FedFM approaches for model training, aggregation, trustworthiness, and incentivization. Key challenges, including how to enable FL to deal with high complexity of computational demands, privacy considerations, contribution evaluation, and communication efficiency, are thoroughly discussed. Moreover, this paper explores the intricate challenges of communication, scalability and security inherent in training/fine-tuning FMs via FL. It highlights the potential of quantum computing to revolutionize the processes of training, inference, optimization and security. This survey also introduces the implementation requirement of FedFM and some practical FedFM applications. It highlights lessons learned with a clear understanding of our findings for FedFM. Finally, this survey not only provides insights into the current state and challenges of FedFM, but also offers a blueprint for future research directions, emphasizing the need for developing trustworthy solutions. It serves as a foundational guide for researchers and practitioners interested in contributing to this interdisciplinary and rapidly advancing field.
@@
Federated learning, foundation models, federated foundation models, large language models, efficient training/aggregation, trustworthiness, incentivization, evaluation, quantum computing.
@@ I Introduction
I-A Motivation
The Artificial Intelligence (AI) landscape is undergoing significant advancement propelled by the emergence of Foundation Models (FMs) [1, 2]. FMs are large-scale machine learning (ML) models that serve as a foundational backbone for a wide range of learning tasks across diverse modalities[3]. They are typically pre-trained on extensive datasets using self-supervised learning methods, which equip them with general and adaptable knowledge for various tasks. While technically representing an evolutionary rather than revolutionary development in AI [2], FMs have significantly reshaped model training and deployment practices. This practice arises from the high costs associated with FM training and the relatively limited scale of available datasets, making the traditional training-from-scratch paradigm impractical.
Recently, FMs with their diverse designs and learning processes have achieved remarkable success in tackling complex tasks that were once considered too difficult or even impossible. Particularly in Natural Language Processing (NLP), FMs such as those in the GPT series [4, 5, 6], LLaMa [7] and PaLM [8] have demonstrated extraordinary capabilities. In computer vision, Segment Anything [9] has emerged as a standout performer. In the realm of Generative AI models, Stable Diffusion [10] has garnered significant attention. Notably, DALL-E [11] and CLIP [12] have demonstrated outstanding performance in multi-modal tasks, bridging the understanding of textual and visual data.
However, the success of FMs comes at a high cost. The enhanced capabilities require a substantial amount of high-quality training data and computational resources. In addition, centralized training of FMs raises concerns about data privacy and security. These challenges underscore the potential value of a collaborative FM training paradigm that enables the distribution of both data resources and computational loads, while preserving data privacy.
Federated Learning (FL) [13] presents a promising solution to these challenges. It is a collaborative learning approach that enhances privacy, leverages distributed datasets, and distributes computational loads. Training and fine-tuning FMs via FL has several advantages. FL can broaden the data horizons for FMs by aggregating diverse data sources from both the public and private sectors [14], including research institutions and industries. It can also distribute computational loads among multiple participants, leading to more efficient resource utilization and mitigating the risk of AI monopolies by major technology companies. On the other hand, the robustness of FMs strengthens the effectiveness of FL, particularly in managing non-iid (non-independent and identically distributed) data. Moreover, the flexibility of FMs in adapting to various downstream tasks facilitates easier personalization of FL models. These benefits highlight the promising potential of integrating FMs with FL.
@@ I-B Related Works and Contributions
Figure 1: Organization of the FedFM survey.
I-C Organization of This Survey Paper
The survey paper is organized as follows. Section II introduces FMs and FL, concepts relevant to our discussion. Section III describes the FedFM taxonomy, which encompasses several key aspects: training and aggregation methods, ensuring trustworthiness, creating effective incentive mechanisms, and managing evaluation. Section IV focuses on aggregation, communication and computational efficiency, highlighting the nuanced considerations required for the effective deployment of FedFM. Section V addresses the critical areas of robustness, privacy and intellectual property for FedFM. Section VI highlights the complexity of developing effective incentive mechanisms for FedFM.
Then, Section VII describes the practical implementation and applications of FedFM in terms of platforms and libraries and hardware consideration. Section VIII summarizes the key insights of three categories of FedFM and connects these lessons to Section IX. Section IX discusses open research challenges of FedFM in terms of memory, computing and communication, while also delving into potential future directions involving quantum computing techniques for FedFM and effective FedFM evaluation. Finally, Section X concludes this FedFM survey paper.
II Background
This section aims to provide an overview of FMs and FL, outlining the fundamental concepts and historical development trajectories that have propelled their current prominence in the field. There are more recent surveys on FMs and FL fields, respectively.
Recent surveys on FMs have highlighted the advancements and applications of these models in various AI tasks. Bommasani et al. [2] provide a detailed analysis of the opportunities and risks associated with FMs, discussing their impact on different AI fields. Another survey by Myers et al. [22] explores the design, training techniques, and deployment challenges of FMs, offering insights into their scalability and performance. Furthermore, Brown et al. [5] review the progress and future trends in FMs, with a particular focus on language models like GPT-3/GPT-4 and their implications for AI research.
Recent surveys have extensively discussed relevant schemes for efficiency, privacy, and other challenges in FL. For instance, Li et al. [23] provide a comprehensive review of FL, covering various optimization techniques, privacy-preserving methods, and system challenges. Kairouz et al. [13] delve into the advancements in FL, emphasizing theoretical foundations, communication efficiency, and security measures. In addition, Yang et al. [24] discuss the applications of FL across different domains, highlighting the practical considerations and future directions for FL research.
@@ II-A Foundation Models
The term FMs was coined in [2] to denote the role of such models as a foundational base, from which numerous task-specific models can be built through adaptation. FMs emphasize the stability of their architecture during training and the consistency of parameters throughout the adaptation phase. This term also signifies a profound shift in AI research and deployment, which began in the field of NLP, where the synergy of large-scale models and transfer learning [25] resulted in the development of powerful models.
The need for FMs arises from several critical challenges and driving forces in the field of AI:
  • High Costs of Training: Training large-scale models from scratch is resource-intensive, requiring substantial computational power and vast amounts of data. FMs address this by being pre-trained on large datasets, enabling them to learn generalized knowledge that can be fine-tuned for specific tasks, reducing the need for extensive resources for each new task. For example, GPT-2 [4] consists of 1.3 billion parameters with a memory footprint of approximately 2.6 GB, while GPT-3 [5] consists of 175 billion parameters with a memory footprint of around 350 GB in half-precision format. GPT-4 [6] has reached the scale of over one trillion parameters.
  • Data Scarcity: Many AI tasks suffer from a lack of sufficient high-quality labeled data. FMs, pre-trained on diverse and extensive datasets, can be adapted to new tasks with limited data, leveraging their pre-existing knowledge to achieve high performance even with smaller datasets. GPT-3, for instance, is trained on a dataset with 300 billion tokens [5], equivalent to hundreds of gigabytes of data. Research suggests a potential depletion of high-quality language data by 2026 [26], raising concerns that data availability might eventually impede the advancement of FMs.
  • Scalability: Traditional models often struggle with scalability when dealing with diverse tasks and large datasets. FMs, with their scalable architectures, can handle a wide range of tasks across different domains, making them versatile and efficient for various applications [27].
  • Transfer knowledge Ability: The ability to transfer knowledge from one task to another is a significant advantage of FMs. By pre-training on broad data and fine-tuning for specific tasks, FMs facilitate transfer learning, allowing models to be quickly adapted to new tasks without starting from scratch [28].
  • Efficiency and Performance: FMs have demonstrated remarkable success in improving efficiency and performance across numerous AI applications. In NLP, for instance, models like GPT series [4, 5, 6], LLaMa[7], and PaLM have set new benchmarks for language understanding and generation. In computer vision, models like Segment Anything and generative models like Stable Diffusion and DALL-E have achieved impressive results. For instance, the LLaMA model requires 2,048 NVIDIA A100 GPUs for a duration of 21 days for training[7], while the vision transformer in CLIP requires 8 core TPUv3 for approximately 30 days for training [29].
  • Unified Architectures: FMs provide a unified architecture that can be applied to multiple tasks, simplifying the model development process and fostering consistency in performance [30]. This unified approach contrasts with the need to develop and train separate models for each task, streamlining the workflow and enhancing efficiency.
Besides, Table I also contrasts FedFM and traditional FMs across a diverse spectrum of topics ranging from data efficiency, data privacy, model performance, system operation, scalability, deployment, consistency and latency. All of these factors have collectively propelled the evolution and widespread adoption of trustworthy AI.
@@ II-B Federated Learning
FL aims to collaboratively train models across multiple decentralized data owners holding potentially sensitive local data in a privacy-preserving manner [13]. The essence of FL lies in its ability to learn a shared model by aggregating locally computed updates, rather than directly accessing or sharing the raw data. This ML paradigm not only enhances privacy and security, but also enables the utilization of distributedly owned data for model training, making it particularly suitable for mission critical applications, such as healthcare [31] and finance [32], where data privacy is of paramount concern. The motivation behind FL is driven by several key factors:
  • Data Privacy and Security: In many domains, especially healthcare and finance, data are highly sensitive and cannot be shared directly due to privacy concerns and regulatory constraints [23]. FL addresses this by ensuring that raw data remain locally and only model updates are shared, significantly reducing the risk of privacy breaches.
  • Utilization of Distributed Data: Organizations often have valuable data that is distributed across various locations. Traditional centralized training methods cannot effectively leverage this distributed data without transferring it to a central server, which is often impractical [24]. FL enables the use of this distributed data for training robust models without requiring data centralization.
  • Reduction of Communication Overhead: Transferring large datasets to a central server for training can be resource-intensive and time-consuming [13]. FL reduces this overhead by only transmitting model updates, which are typically much smaller in size compared to the raw data.
  • Scalability and Collaboration: FL facilitates collaboration among multiple entities, such as hospitals or financial institutions, allowing them to build more accurate and generalizable models by leveraging diverse datasets. This collaborative approach enhances the scalability of AI solutions across different domains[24].
  • Enhanced Model Robustness: By training on data from multiple sources, FL can produce models that are more robust and generalizable [23]. This is particularly important for applications where the model needs to perform well across different environments and conditions.
@@ III-A Efficient Federated Training and Aggregation: The Foundation of FedFM
This part of the review underscores the evolving landscape on FedFM research, detailing the challenges, strategies, and innovations in this interdisciplinary field. The focus on scaling, communication, and computational efficiency reflects the nuanced considerations required to effectively deploy FedFM.
@@ III-B Trustworthiness: A Crucial Pillar of FedFM Integrity
Trustworthy FedFM address the critical areas of robustness, privacy preservation, and intellectual property in the context of FedFM, exploring the challenges and proposing solutions to ensure the integrity and security of these systems. We summarize key trends in each of these areas:
  • Robustness: Research in this area discusses how poisoning attacks, both untargeted and targeted, aim to compromise the integrity of global models in FL systems. Untargeted attacks disrupt the training process to prevent convergence, while targeted attacks manipulate the model output subtly. The complexity and heterogeneity of FedFM training tasks make these attacks particularly challenging to execute and defend against. Byzantine-robust aggregation rules are also being proposed to counteract poisoning attacks, including geometrical outlier detection, top performance selection, and other hybrid schemes. However, the effectiveness of these defenses has been questioned in FedFM settings, calling for novel approaches tailored to its complexities.
@@
  • Intellectual Property Protection: Research in this area emphasizes the importance of intellectual property protection for safeguarding the ownership of FedFM against unauthorized usage, such as model theft. Two primary strategies are being explored in the context of FedFM: black-box fine-tuning and watermarking. Black-box fine-tuning allows for model adaptation while keeping the core intellectual property intact, whereas watermarking embeds identifiable markers within the model to assert ownership. The integration of these methods ensures that FedFM can be securely and efficiently utilized across distributed networks. Nevertheless, the effectiveness of these strategies in the diverse and complex environments typical of FedFM has prompted ongoing discussions and investigations into more robust and tailored approaches.
This part of the review on FedFM research underscores the intricate balance required to achieve trustworthiness, emphasizing the need for innovative solutions to address the dual challenges of robustness against poisoning attacks, privacy preservation, and intellectual property. The complexity of FMs and the FL environment necessitates a re-evaluation of traditional defense mechanisms and development of new strategies to ensure the integrity and privacy of such advanced learning systems.
@@ IV Towards Efficient FedFM
In this section, we review and discuss current literature on the topic of efficient collaborative model training/fine-tuning and aggregation for FedFM, focusing on emerging settings and scenarios that arise as a result of combining FL with FMs. We further break down each dimension into sub-areas for more detailed analysis (Fig. 3).
@@
14:        $\bm{\theta}_{i}^{t}\leftarrow\bm{\theta}^{t}-\eta\nabla\mathcal{L}(\bm{\theta}^{t};b)$
15:        // Q2) How to improve the computational efficiency so that FMs can be efficiently hosted and updated on FL clients?
@@
18:  return $\bm{\theta}_{i}^{t}$ to the server
19:  // Q3) How to improve communication efficiency when $\bm{\theta}_{i}^{t}$ is large?
20:
21:  procedure Aggregate($\{\bm{\theta}_{i}^{t+1}\}_{i=1}^{m}$)
22:  FedAvg [33]: $\bm{\theta}^{t+1}\leftarrow\frac{1}{m}\sum_{i=1}^{m}\bm{\theta}_{i}^{t+1}$
23:  or
24:  // Q1) Aggregation methods for FMs to enhance performance?
@@ IV-A FedFM Training and Aggregation
To grasp the challenges involved in integrating FMs into the FL process, a revision of the general FL procedure is necessary. FL typically operates under the control and decision-making of the central server with the primary purpose of collaboratively training a Deep Learning (DL) model $\bm{\theta}$ across $N$ clients over multiple rounds, as outlined in Algorithm 1. The integration of FMs, with their considerable sizes, introduces challenges at various stages of the FL process. These challenges can be distilled into three critical research questions (highlighted in yellow in Algorithm 1):
  • Q1)
@@
Figure 4: Illustration of efficient FedFM.
Therefore, our survey examines innovative methods and strategies useful for FedFM from three dimensions: 1) aggregation methods for enhanced effectiveness, 2) techniques for boosting computational efficiency, and 3) techniques for improving communication efficiency. We further break down each domain into specific areas as shown in Fig. 4.
@@ IV-B Scaling Efforts for FedFM
TABLE III: Sizes of FMs and Aggregation under FL Setting
Methods Year $\downarrow$ Size (M) Aggregation Model Description
FedPAQ [36] 2020 0.2 FedSGD Implement quantization techniques to reduce communication overhead in FL by compressing model updates
HeteroFL [37] 2020 11 FedAvg Support heterogeneous model architectures across different clients, optimizing the training process based on individual device capabilities
LotteryFL [38] 2020 138 FedAvg Apply the lottery ticket hypothesis to identify and retain efficient sub-networks to improve training efficiency
FjORD [39] 2021 11 FedAvg Utilize ordered dropout to maintain a structured hierarchy of knowledge within the model, facilitating efficient extraction of sub-models
H-FL [40] 2021 138 FedAvg Address statistical heterogeneity among clients by employing lossy SVD and bias correction mechanisms
PruneFL [41] 2022 132 FedAvg Introduce a two-stage pruning process to reduce model size while preserving performance, tailored for FL environments
FedPM [42] 2022 12 FedAvg Use a binary mask approach inspired by the lottery ticket hypothesis to prune models, enhancing efficiency
FedTiny [43] 2022 132 FedAvg Enhance adaptability by using batch normalization statistics for initializing models, suitable for diverse data distributions across clients
SoteriaFL [44] 2022 0.05 FedSGD Combine differential privacy and communication compression to protect data while reducing the overhead in FL
FedPrompt [45] 2022 223 FedAvg Focus on prompt tuning techniques for fine-tuning pre-trained models in federated environments, optimizing performance on specific tasks
FedBERT [46] 2022 117 FedAvg Adapt the BERT model for FL, enabling pre-training and fine-tuning across distributed datasets
FedCLIP [47] 2023 85 FedAvg Adapt the CLIP for federated settings by leveraging adapter- based fine-tuning to handle multi-modal tasks effectively
FedPEFT [48] 2023 85 FedAvg Introduce parameter-efficient fine-tuning methods that modify only few parameters, reducing the burden on client devices
FedPETuning [49] 2023 125 FedAvg Conduct benchmark analyses of adapter-based fine-tuning techniques, assessing their efficiency in FL
SLoRA [50] 2023 67 FedAvg Optimize LoRA for non-IID federated settings, using parameter moderation to handle diverse data distributions
FedOBD [51] 2023 17 FedAvg Segment FMs into semantic blocks for selective transmission, reducing communication costs while maintaining performance
FedIT [52] 2023 7,000 FedAvg Apply instruction tuning to the LLaMA model, allowing it to handle a variety of tasks simultaneously in a FL setting
FwdLLM [53] 2024 7,000 FedSGD Combines Backpropagation-free training with parameter-efficient methods to adapt FMs for mobile devices with limited resources
We list the methods reviewed alongside with the largest base models involved in their evaluations in Table III and Table IV. The substantial difference for model sizes between existing FL research and centralized FMs might cast doubts on the feasibility and practicality of the FL techniques for FedFM, calling for further experimental studies involving large-scale FMs. Such an effort depends on the availability of frameworks that can streamline implementation. Thus, in Section VII-A, we analyze the support for FMs in existing FL frameworks.
@@
Methods Year $\downarrow$ Size (M) Model Description
BERT [54] 2018 340 A transformer-based model pre-trained on large text corpora, widely used for various NLP tasks
GPT-1 [55] 2018 117 The first Generative Pre-trained Transformer (GPT) model, designed for language modeling tasks, setting the foundation for subsequent versions
GPT-2 [4] 2019 1,500 An improved version of GPT with significantly more parameters, enhancing performance on a range of NLP tasks
GPT-3 [5] 2020 175,000 A large-scale language model with 175 billion parameters, capable of performing a wide variety of language generation and comprehension tasks
ViT [56] 2020 632 The Vision Transformer model by applying transformer architecture to image classification tasks, demonstrating high performance on standard benchmarks
T5 [57] 2020 11,000 The Text-to-Text Transfer Transformer, designed to convert all NLP tasks into the text-to-text format, achieving state-of-the-art performance
DALL-E [58] 2021 12,000 An image generation model that creates images from textual descriptions, showcasing the capabilities of generative models
CLIP [59] 2021 1,000 A multi-modal model that can understand and generate images and text, trained to connect textual descriptions with images
Gato [60] 2022 1,200 A generalist agent by DeepMind capable of performing various tasks across different modalities using a single neural network
PaLM [61] 2023 540,000 A Pathways Language Model designed for scaling up to very large models, demonstrating strong performance across a wide range of NLP tasks
LLaMA [62] 2023 69,000 A large-scale language model designed for multi-task learning, integrating capabilities across different domains to enhance versatility
GPT-3.5 [63] 2023 20,000 GPT-3.5 Turbo models can understand and generate natural language or code and have been optimized for chat but work well for non-chat tasks as well
[…]
Chinchilla [64] 2023 70,000 An efficient language model designed by DeepMind, focusing on optimized training efficiency and resource utilization
Flamingo [65] 2023 80,000 A model designed for visual understanding tasks, integrating both vision and language inputs for improved performance
LLAVA [66] 2024 15,000 A vision-language model aimed at handling multi-modal tasks, combining visual and textual data to improve task performance
LLaMA 3 [67] 2024 70,000 The models are pre-trained on 15 trillion tokens of text gathered from publicly available sources with the instruct models fine-tuned on instruction datasets
GPT-4o [68] 2024 Not disclosed GPT-4o (“o” for “omni”) has the same high intelligence as GPT-4 Turbo but is much more efficient—it generates text 2x faster and is half cheaper
LLaMA 3.1 [69] 2024 405,000 An auto-regressive LLM that uses an optimized transformer architecture. The tuned versions use supervised fine-tuning and reinforcement learning with human feedback to align with human preferences for helpfulness and safety